Yellow​ Press, Inc., buys paper in​ 1,500-pound rolls for printing. Annual demand is 2 comma 250 rolls. The cost per roll is ​$1 comma 000 ​, and the annual holding cost is 20 percent of the cost. Each order costs ​$35 . a. How many rolls should Yellow Press order at a​ time? Yellow Press should order nothing rolls at a time. ​(Enter your response rounded to the nearest whole​ number.)
Answer #1

EOQ = SQRT(2 * DEMAND * ORDERING COST / HOLDING COST)

EOQ = SQRT(2 * 2250 * 35 / 200) = 28.06

Yellow press should order 28 rolls at a time.

Do give a positive rating if you think this answer helped